John bought two horses at rs 20,000 each.he sold one horse at 15% gain. But he had to sell the second horse at a loss.if he had suffered a loss of rs 1800 on the whole transaction,find the s.p of the second horse.
Answer by mananth(16946)


Profit on first horse = 20,000*15%
15/100 * 20000= Rs. 3000
SP = 23000
Loss on the second horse = x
SP= CP-loss
SP= 20000-x
SP of both = 23000+20000-x
CP of both = 40000
CP-SP=Loss = 1800
40000-(43000-x)= 1800
40000-43000+x=1800
-3000+x=1800
x= 4800
SP= cp-loss
SP= 20000-4800
SP= 15200
He sold second horse for Rs. 15200. 00
m.ananth@hotmail.ca